give in :angle ABC ac=4x,bc=6x-6,ab=5x-8 and angle a= angle b . find the perimeter of angle abc

Answer by mananth(16946)

Triangle ABC ac=4x,bc=6x-6,ab=5x-8 and angle a= angle b
angle a = angle b
side opp to angle a = side opposite to angle b
BC= AC
6x-6 = 4x
6x-4x=6
2x=6
x=3
...
AC = 4x = 12
bc=6x-6 = 12
ab = 5x-8 = 7
...
Add up to get the perimeter = 31